### v1

# Retrieval of fundamental material parameters of monolayer transition metal dichalcogenides from experimental exciton energies: An analytical approach

Abstract
We propose a straightforward and highly accurate method for determining material parameters such as screening length, bandgap energy, exciton reduced mass, and dielectric constant of the surrounding medium using experimental magnetoexciton energies available for monolayer transition metal dichalcogenides (TMDCs). Our approach is based on analytical formulations that allow us to calculate the screening length $r_{0}$ and bandgap energy $E_{g}$ directly from the experimental $s$-state exciton energies $E_{1s}$, $E_{2s}$, and $E_{3s}$. We also establish a relationship between the surrounding dielectric constant $\kappa$ and the exciton reduced mass $\mu$. This relationship simplifies the Schrödinger equation for a magnetoexciton in a TMDC monolayer, transforming it into a one-parameter equation that depends only on a single material parameter, $\mu$. Therefore, the fitting method is easy to implement, as it compares calculated magnetoexciton energies with experimental data to determine the true exciton reduced mass, the best-fit value of $\mu$. By applying this method, we extract material parameters $E_{g}$, $r_{0}$, $\mu$, and $\kappa$ from the magnetoexciton energies of monolayer TMDCs (WSe2, WS2, MoSe2, and MoS2) encapsulated by hexagonal boron nitride (hBN) slabs in various current experiments. Additionally, we provide key material properties such as diamagnetic coefficients and exciton radii, which are calculated using the extracted parameters.
Exciton, transition metal dichalcogenides, material parameters, retrieval, analytical exciton energy
I Introduction
Two-dimensional semiconductors with thickness of atomic scales, such as monolayer transition metal dichalcogenides (TMDCs), have attracted significant attention over the past decade because they enable the fabrication of ultra-thin electronic devices with exceptional performance [1, 2, 3, 4]. High promises of these materials come from their incredible properties, which include bandgap tunability, high optical sensitivity, strong spin-orbit coupling, valley polarization, considerable excitonic binding energy, and the high thermal stability of exciton at room temperature [5, 6, 7, 8, 9, 10]. Therefore, a precise knowledge of the intrinsic optoelectric properties of these TMDCs, like their energy bandgap $E_{g}$, exciton reduced mass $\mu$, screening length $r_{0}$ (related to the two-dimensional (2D) static polarizability $\chi_{2D}$) as well as the effective dielectric constant $\kappa$ of the surrounding medium is essential for advancing the development of TMDC–based devices.
There are various methods available for determining the optoelectric properties of TMDCs. For instance, the exciton reduced mass can be directly determined via effective electron and hole masses extracted from angle-resolved photoemission spectroscopy (ARPES), which can experimentally detect energy versus momentum maps [11, 12, 13, 14]. Meanwhile, the bandgap energy can be probed by time-resolved ARPES [13], scanning tunneling microscopy (STM) [15, 16, 17], or photoluminescence excitation (PLE) spectra [18, 19]. From the theoretical side, ab initio calculations based on density functional theory (DFT), time-dependent DFT, or GW-Bethe-Salpeter-equation (GW-BSE) could give some predictions about the exciton reduced mass, bandgap energy, or even 2D polarizability [20, 21, 22, 23, 24, 25]. Unfortunately, balancing accuracy and cost among these experimental measurements or theoretical calculations is challenging.
Nevertheless, exciton spectroscopy provides other ways to probe intrinsic optoelectric properties of materials, combining theoretical knowledge of exciton spectra and experimental measurements via optical spectroscopies. Recent studies [26, 27, 28, 29] revealed the exciton reduced mass $\mu$ from the optical spectroscopy of magnetoexcitons at a high magnetic field (65 to 91 tesla) in the same manner as people did in 3D perovskite semiconductors [30, 31]; however, such a magnetic field is not theoretically large enough for their approximation (see a discussion in Ref.  [32]). Another promising approach is to use both PLE and two-photon measurements as people did in 3D semiconductors [31]. This approach requires a theoretical model to describe the relationship between the exciton spectra of $s$ and $p$ states, which is currently unavailable due to the nonhydrogenic nature of exciton spectra in TMDC monolayers. Indeed, in this case, unlike 3D semiconductors, the dimensionality reduction causes a screening effect in the Coulomb interaction between electron and hole [21, 33], which makes them coupled by the Rytova-Keldysh potential [34]. Thus, exciton spectra in TMDC monolayers are much more complicated than the well-known 2D hydrogenic model [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45].
Figure 1: Schematic flowchart of extracting the material parameters (the exciton reduced mass $\mu$, screening length $r_{0}$, surrounding dielectric constant $\kappa$, and bandgap energy $E_{g}$) from experimental exciton energies.
In the spirit of the availability of magnetoexciton spectroscopy for monolayer TMDCs, our group developed a method of probing material properties, $E_{g}$, $\mu$, $r_{0}$, and $\kappa$, from the magnetoexciton spectra [46, 47, 32]. Particularly, in Ref. [32], the exact numerical calculations of exciton energy spectra via the so-called Feranchuk-Komarov (FK) operator method [48, 49] allowed us to examine the sensitivity of magnetoexciton energies on the material parameters from which we established an efficient fitting scheme for probing material properties. Notwithstanding its accuracy, our numerical fitting scheme relied purely on extensive numerical calculations, preventing us from exploring the underlying physics of the method.
Learning from the case of 3D semiconductors [30, 31], we believe that an analytical model for exciton energy spectra at a zero field can simplify our retrieval process and make sense of the retrieved properties. Fortunately, an analytical expression for s-state exciton energy spectra at zero fields was proposed within the regulated perturbation theory [45]. In the present study, such an analytical formula allows us to develop a formulation for bandgap energy, screening length, and surrounding dielectric constant directly from the zero-field experimental exciton energies. These formulas also enormously reduce the numerical fitting scheme for the exciton reduced mass via magnetoexciton spectra. Figure 1 illustrates our proposal to retrieve material parameters from experimental magnetoexciton energies.
The rest of the present paper is as follows. In Section II, we derive formulas for the bandgap energy $E_{g}$, screening length $r_{0}$, and surrounding dielectric constant $\kappa$. The bandgap energy $E_{g}$ and screening length $r_{0}$ can be calculated directly from zero-field s-state exciton energies. Meanwhile, the formula for the surrounding dielectric constant $\kappa$ also contains the exciton reduced mass $\mu$. Section III is dedicated to the numerical fitting scheme for the exciton reduced mass from the magnetoexciton energies. Section IV provides material parameters $E_{g}$, $r_{0}$, $\mu$, and $\kappa$ retrieved from the experimental exciton energies available for monolayer TMDCs by our method and compares them with the available data. Some other parameters, such as the diamagnetic constants and exciton radii calculated from the retrieved fundamental parameters, are also provided in this section. Section V presents our conclusions.
II Analytical formulas for the bandgap energy, screening length, and surrounding dielectric constant
Based on the Rytova-Keldysh potential describing the hole-electron interaction in monolayer TMDCs, Dinh et al. [45] proposed an analytical formula for the $s$-state exciton binding energies. For use in the present study, we rewrite it as
$\varepsilon_{ns}=-\eta\frac{P_{n}(\xi)}{\left(n-0.5+\xi\right)^{2}}\,\textrm{Ry},\quad n=1,2,3,\dots$ (1)
where Ry $=13605.69$ meV is a Rydberg unit of energy; $P_{n}(\xi)={\textrm{exp}}(a_{n}\Delta)/(1+b_{n}\Delta)$ with $\Delta=0.225/\xi-0.525$, $a_{n}=4.3/n-6.6/n^{2}+3.2/n^{3}$, and $b_{n}=5.3/n-7.6/n^{2}+4.0/n^{3}$. Here, the dimensionless parameter $\eta$ is the rescaling factor of Rydbergy energy, while the dimensionless parameter $\xi$ captures the screening effect from dimensionality reduction. These parameters are defined as
$\eta=\frac{\mu}{\kappa^{2}\,m_{e}},\quad\xi=0.479\,\frac{r_{0}}{a_{0}}\frac{\mu}{\kappa^{2}\,m_{e}},$ (2)
where $\mu$ is the exciton reduced mass; $m_{e}$ is the electron mass; $\kappa$ is the dielectric constant of the surrounding medium; $r_{0}$ is a screening length; $a_{0}=0.0529$ nm is the Bohr radius. We note that formula (1) works well only in the range of $0.290.85>\xi>$ 0.29.
On the other hand, in the experimental photoabsorption spectrum, exciton peaks are located at
$E_{ns}=\varepsilon_{ns}+E_{g},$ (3)
where $E_{g}$ is a bandgap energy of the TMDC monolayer, which needs to be retrieved. From Eqs. (1) and (3), we can inversely define the bandgap energy $E_{g}$ and parameters $\eta$, $\xi$ from the experimental exciton energies $E_{1s}$, $E_{2s}$, and $E_{3s}$ since there are three equations versus three unknowns, as shown by the following scheme:
${\textrm{Experimental}}\,(E_{1s},E_{2s},E_{3s})\rightarrow(\eta,\xi,E_{g}).$ (4)
Indeed, from Eqs. (1) and (3), we can find out that the following ratio
$\delta(\xi)$ $=$ $\frac{\varepsilon_{3s}-\varepsilon_{2s}}{\varepsilon_{2s}-\varepsilon_{1s}}=\frac{E_{3s}-E_{2s}}{E_{2s}-E_{1s}}$ (5)
$=$ $\frac{(1+2\xi)^{2}}{(5+2\xi)^{2}}\,\frac{(3+2\xi)^{2}P_{3}(\xi)-(5+2\xi)^{2}P_{2}(\xi)}{(1+2\xi)^{2}P_{2}(\xi)-(3+2\xi)^{2}P_{1}(\xi)}$
is a function of only the screening parameter $\xi$ and does not depend on the Rydberg rescaling parameter $\eta$. Therefore, the experimental data of $E_{1s}$, $E_{2s}$, and $E_{3s}$ for different materials such as WSe2, WS2, MoSe2, and MoS2 [27, 29, 37, 38, 28, 36] yield us to the experimental quantity $\delta_{exp}={(E_{3s}-E_{2s})}/(E_{2s}-E_{1s})$, from which we can determine the screening parameter $\xi$ by solving the following equation:
$\delta(\xi)=\delta_{{exp}}\,.$ (6)
For monolayer TMDCs on the hexagonal boron nitride (hBN) substrate, as studied in the most recent references [27, 29, 37, 38, 28, 36], the quantity $\xi$ lies in the working range of formula (1) for exciton binding energies. Since the function $\delta(\xi)$ is single-valued and monotonically increases as $\xi$ increases in its range, as demonstrated in Fig. 2, Eq. (6) has a unique solution. This kind of equation is easy to solve numerically; however, we approximately solve it to obtain an analytical solution as:
$\xi=1.1219-3.821\,\sqrt{0.2036-\delta_{exp}}\,.$ (7)
Figure 2: Ratio $\delta(\xi)$ by formula (5) as a function of variable $\xi$ (black line) plotted with some experimental data [27, 28, 36] of $\delta_{\text{exp}}=(E_{2s}-E_{1s})/(E_{3s}-E_{2s})$ marked by different symbols. Also solution $\xi(\delta)$ (7) plotted for comparison (red dots).
After determining the screening parameter $\xi$ from Eq. (6), we can directly obtain the Rydberg rescaling parameter $\eta$ as well as the bandgap energy $E_{g}$ by solving Eq. (3) for the first two states (n=1 and n=2). The results are
$\eta=\frac{(1+2\xi)^{2}(3+2\xi)^{2}}{(3+2\xi)^{2}P_{1}(\xi)-(1+2\xi)^{2}P_{2}(\xi)}\frac{E_{2s}-E_{1s}}{4\textrm{Ry}},$ (8)
$E_{g}=E_{1s}-\frac{(3+2\xi)^{2}P_{1}(\xi)\left(E_{2s}-E_{1s}\right)}{(1+2\xi)^{2}P_{2}(\xi)-(3+2\xi)^{2}P_{1}(\xi)}.$ (9)
Using Eqs. (8) and (9), we can calculate $\eta$ and $E_{g}$ from the experimental exciton energies ($E_{1s}$, $E_{2s}$) and the value $\xi$, which is in turn determined from the experimental exciton energies $E_{1s}$, $E_{2s}$, and $E_{3s}$ by Eq. (6).
Knowing the parameters $\eta$ and $\xi$, we can directly achieve the screening length from Eqs. (2) as
$r_{0}=2.088\,\frac{\xi}{\eta}\,a_{0}\,.$ (10)
Also, the surrounding dielectric constant $\kappa$ can be obtained from the first equation in Eqs. (2), which leads to the formula
$\kappa=\sqrt{\frac{\mu}{\eta\,m_{e}}}\,.$ (11)
Equation (11) does not mean the surrounding dielectric constant $\kappa$ is dependent on the exciton reduced mass $\mu$ but serves for our retrieval scheme only.
For this Section, we conclude that the bandgap energy $E_{g}$ and screening length $r_{0}$ can be determined from the experimental exciton energies $E_{1s}$, $E_{2s}$, and $E_{3s}$ by Eqs. (9) and (10), independently of knowing the values of the surrounding dielectric constant $\kappa$ and the exciton reduced mass $\mu$. This conclusion also explains the numerical results of Ly et al. [32], where $r_{0}$ remains unchanged even when varying the values of $\kappa$ and $\mu$. Besides, the relationship (11) allows reducing one more material parameter, saying the surrounding dielectric constant $\kappa$, from the Schrödinger equation. These three formulas, Eqs. (9), (10), and (11), are essential because they remove the need to determine all structural parameters ($\mu$, $\kappa$, $r_{0}$, and $E_{g}$) simultaneously, remaining to achieve only the exciton reduced mass $\mu$.
III Exciton reduced mass from magnetoexciton energies
To determine the exciton reduced mass $\mu$, we suggest a method based on fitting the calculated magnetoexciton energies with the experimental data.
For the fitting method, we need to solve the following Schrödinger equation for the s-state exciton in monolayer TMDCs as
$\left\{-2\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)+\frac{1}{8}\frac{\mu^{2}}{m_{e}^{2}}B^{2}(x^{2}+y^{2})\right.\qquad$
$\left.-\frac{2}{\kappa}\int\limits_{0}^{+\infty}\frac{dq}{\sqrt{1+\alpha^{2}q^{2}}}\;\textrm{e}^{-qr}-\varepsilon\right\}\psi(x,y)=0,\quad$ (12)
where $\alpha=r_{0}/(\kappa a_{0})$; $r=\sqrt{x^{2}+y^{2}}$; energy $\varepsilon$ and coordinates $x,y$ are given in the Rydberg unit Ry and Bohr radius $a_{0}$, respectively; B is dimensionless magnetic intensity measured in the unit of $B_{0}=(m_{e}/\hbar e)\text{Ry}$; $\hbar$ and $e$ are the reduced Planck constant and electron electric charge, respectively. Here, in Eq. (III), we use the Laplace transform for the Rytova-Keldysh potential describing the electron-hole interaction, given in Ref.  [32].
The Schrödinger equation (III) contains three material parameters: $\kappa$, $r_{0}$, and $\mu$. However, we can exclude two of them, $\kappa$ and $r_{0}$, from the equation using formulas (10) and (11). Consequently, the equation consists of only one material parameter, $\mu$. Previous studies show the strong dependence of the magnetoexciton energies on the exciton reduced mass. Therefore, we consider $\mu$ a variable parameter and vary it so that the calculated magnetoexciton energies $\varepsilon_{ns}(B),\,(n=1,2,3,4)$ best fit with the experimental data $E_{ns}(B)-E_{g}$, where the bandgap energy $E_{g}$, in turn, can be calculated by formula (9) from the experimental data. This way, we retrieve the exciton reduced mass $\mu$ from the experimental magnetoexciton energies $E_{ns}(B)$. We note that a numerical method based on the Feranchuk-Komarov method was proposed in Ref. [32] to solve the Schrödinger like Eq. (III) with the codes available upon request. This method is speedy and gives highly accurate magnetoexciton energies.
For demonstration purposes, in Fig. 3 (a), we plot the magnetoexciton energy spectrum for monolayer WSe2 on the hBN substrate calculated with different values of $\mu$ (0.17, 0.18, and 0.19 $m_{e}$). Other material parameters ($E_{g}$, $r_{0}$, $\kappa$) are obtained by Eqs. (9), (10), and (11) with the experimental exciton energies ($E_{1s}$, $E_{2s}$, $E_{3s}$) of Ref. [29]. We also plot the experimental magnetoexciton energies from Ref. [29] for comparison. As discussed above, by adjusting $\mu$ so that the theoretical model agrees with experimental data, we obtain the best-fit values of the exciton reduced mass, $\mu=0.18m_{e}$. Other figures, Figs. 3 (b) and (c), are plotted in the same manner but for experimental data of Refs. [27] and [37].
Figure 3: Magnetoexciton energies in monolayer WSe2 encapsulated by hBN slabs, calculated by the numerical method (lines) compared with experimental data (blue triangles) of Refs. [29] (a), [27] (b), and [37] (c). Bandgap energy $E_{g}$ and screening length $r_{0}$ are retrieved from the experimental data of these references. Exciton reduced masses for these cases are the best-fit values (black line), which leads to the surrounding dielectric constants: $\kappa=4.41$ (a), $4.26$ (b), and $3.35$ (c).
IV Fundamental material parameters of monolayer TMDCs from the available experimental magnetoexciton energies
From the above-formulated analytical expressions, we arrive at a practical scheme, shown in Fig. 1, to retrieve material parameters from experimental magnetoexciton energies.
Applying this retrieval scheme to the experimental data of Refs. [27, 29, 37, 36, 38], we retrieve the material parameters for monolayer WSe2. We provide them in Table 2 and compare them with the ones presented in these references. Knowing the material parameters, we also calculate the other essential properties, such as the diamagnetic constants ($\sigma_{1s}$, $\sigma_{2s}$) and exciton radii ($r_{1s}$, $r_{2s}$). More data retrieved in the same way from Ref. [28] for other TMDC monolayers, such as WS2, MoS2, and MoSe2, are also given in Table 2.
The first remark from Table 2 is that although various experiments were performed for the same material, monolayer WSe2 encapsulated by hBN slabs, the surrounding dielectric constants $\kappa$ from these experiments differ due to variations in the experimental setups. However, the exciton reduced mass $\mu$ is weakly dependent on the surrounding dielectric constant $\kappa$, as shown in Table 2. For various experiments, it almost equals $0.19\,m_{e}$. With the exciton reduced mass $\mu$ known, our method is also applicable to photoluminescence experiments without a magnetic field, as shown in Table 2 for Ref. [38]. Indeed, we can determine the bandgap energy $E_{g}$ and screening length $r_{0}$ using Eqs. (9) and (10) within the experimental exciton energies $E_{1s}$, $E_{2s}$, and $E_{3s}$. The surrounding dielectric constant $\kappa$ is then calculated by Eq. (11), putting the exciton reduced mass’s value $\mu=0.19\,m_{e}$, taken from other experiments. Moreover, the surrounding dielectric constants for different experiments are nearly identical (except in Chen et al. [37]), with values approximately equal to the bulk h-BN dielectric constant of 4.5, as referenced in [5].
Table 1: Fundamental optoelectronic material parameters (bandgap energy $E_{g}$, screening length $r_{0}=2\pi\chi_{\text{2D}}$ with 2D static polarizability $\chi_{\text{2D}}$, exciton reduced mass $\mu$, surrounding dielectric constant $\kappa$) for monolayer WSe2 retrieved from magnetoexciton energies of different experiments [27, 29, 37, 36, 38] using analytical expressions (9), (10), and (11) and compared with those given in these references.
WSe2 $r_{0}$ $\mu$ $\kappa$ $E_{g}$ $E_{b}$ $\Delta E_{1s-2s}$ $\Delta E_{2s-3s}$ $\sigma_{1s}$ $r_{1s}$ $\sigma_{2s}$ $r_{2s}$
(nm) $(m_{e})$ (eV) (meV) (meV) (meV) ($\mu$eV/T2) (nm) ($\mu$eV/T2) (nm)
Stier et al. [27] – exp. (169) 130 22 0.31$\pm$0.02 1.7$\pm$0.1 4.6$\pm$0.2 6.6$\pm$0.4
[27] – theory 4.50 0.20 4.50 1.890 167 124 21.3 1.67 6.96
Present work’s theory 4.34 0.19 4.26 1.892 168.6 129.9 22.0 0.33 1.69 5.69 7.01
Liu et al. [29] – exp. (167) 131 21 0.24$\pm$0.1 1.6$\pm$0.4 6.4 $\pm$0.2 $8.24\pm$0.13
[29] – theory 5.00 0.20 3.97 1.884 172 0.31 4.86
Present work’s theory 3.89 0.18 4.41 1.879 167.2 130.9 21.0 0.36 1.71 6.46 7.27
Chen et al. [37] – exp. (177) 131 26 0.5 2.2 5.8 7.6
[37] – theory 4.51 0.22 4.50 1.900 170 128 23 0.25 1.6 4.18 6.5
Present work’s theory 5.52 0.19 3.35 1.907 179.9 130.9 26.0 0.34 1.72 4.89 6.5
Molas et al. [36] – exp. (167) 129 21.5
[36] – theory 4.50 0.20 4.50 1.873 167 128.0 21.2
Present work 4.25 0.19 4.35 1.873 166.8 129.1 21.5 0.33 1.69 5.73 7.04
Takahashi et al. [38] – exp. (167) 129.1 21.6
[38] – theory 3.45 0.20 5.01 1.862 164
Present work 4.28 0.19* 4.33 1.865 167.0 129.1 21.6 0.33 1.69 5.72 7.03
Table 2: Fundamental optoelectronic material parameters (bandgap energy $E_{g}$, screening length $r_{0}=2\pi\chi_{\text{2D}}$ with 2D static polarizability $\chi_{\text{2D}}$, exciton reduced mass $\mu$, surrounding dielectric constant $\kappa$) for monolayer WS2, MoSe2, and MoS2 retrieved from magnetoexciton energies of the experiment [28] using analytical expressions (9), (10), and (11) and compared with those given in this reference.
$r_{0}$ $\mu$ $\kappa$ $E_{g}$ $E_{b}$ $\Delta E_{1s-2s}$ $\Delta E_{2s-3s}$ $\sigma_{1s}$ $r_{1s}$ $\sigma_{2s}$ $r_{2s}$
(nm) $(m_{e})$ (eV) (meV) (meV) (meV) ($\mu$eV/T2) (nm) ($\mu$eV/T2) (nm)
WS2
Goryca et al. [28] – exp. (181) 141 23
[28] – theory 3.40 0.175 4.35 2.238 180 1.8
Present work’s theory 3.74 0.175 4.10 2.239 180.9 140.9 23.0 0.36 1.69 6.33 7.10
MoSe2
Goryca et al. [28] – exp. (232) 171 33
[28] – theory 3.90 0.35 4.40 1.874 231 0.07 1.1
Present work’s theory 4.03 0.355 4.28 1.875 231.6 171.1 32.8 0.08 1.1 1.07 4.15
MoS2
Goryca et al. [28] – exp. (221) 170 29
[28] – theory 3.40 0.27 4.40 2.161 222 0.12 1.2
Present work’s theory 3.34 0.275 4.46 2.160 220.9 170.0 28.9 0.12 1.23 2.06 5.07
The second remark from Tables 2 and 2 is the slight discrepancy between the bandgap energies calculated by our method and those in the references. By applying the bandgap energies we calculated, the experimental binding energies for the $1s$ state (in the parentheses) coincide with the prediction of our theory. The consistency of our predicted energy transitions $\Delta E_{1s-2s}$ and $\Delta E_{2s-3s}$ to the experimental data is much better than the theoretical modeling by the Keldysh potential in Refs. [27, 29, 37] or by the Kratzer potential in Ref. [36] for the electron-hole interaction. We note that the study [29] does not provide the numerical data for the calculated exciton energies. Also, the work [28] employs the fitting method to determine the material parameters. However, the different ways of choosing the bandgap energy cause a slight discrepancy compared with our results for the parameters. The material parameters provided in Tables 2 and 2 are also consistent with our previous study [32] for monolayers WSe2 and WS2.
Although Figs. 3 (a), (b), and (c) validate our retrieved material parameters well for the case of monolayer WSe2, we also plot the magnetoexciton energies calculated with material parameters in Table 2 retrieved from the experimental data of Ref. [28] for monolayers WS2, MoSe2, and MoS2 and compare them with the experimental energies (see Figs. 4 (a), (b), and (c)).
Figure 4: Magnetoexciton energies (black lines) of monolayer WS2 (a), MoSe2 (b), and MoS2 (c), calculated with the retrieved material parameters and compared with the experimental data (blue triangles) in Ref. [28].
Figure 5: Exciton binding energies calculated by numerical method with the retrieved material parameters for monolayer TMDCs in Tables 2 and 2 and compared with the experimental in Refs. [27, 29, 37, 36, 38, 28].
Besides, we confirm the validity of the retrieved material parameters for the free-field case when $B=0$. Putting the complete parameters ($\mu$, $\kappa$, and $r_{0}$) into the Schrödinger equation (III) and solving it using the method given in Ref. [32], we obtain exact numerical exciton energies and compare them with the experimental ones used to retrieve the material parameters. We show the consistency between the theory and experiments in Fig. 5 for monolayer TMDCs used in Tables 2 and 2 [27, 29, 37, 36, 38, 28].
V Conclusion
This study proposes an analytical approach for retrieving fundamental material parameters such as the bandgap energy, screening length, exciton reduced mass, and surrounding dielectric constant from the experimental magnetoexciton energies in monolayer transition metal dichalcogenides (TMDCs). Unlike previous fitting methods, the approach presented here allows for the independent determination of the bandgap energy and screening length by the analytical formulas of the field-free exciton energies for $s$ states, with the fitting procedure is applied for only one parameter, the exciton reduced mass. The surrounding dielectric constant is then also determined through an analytical formula. By applying the suggested method to the currently available experimental data for exciton energies in monolayer TMDCs encapsulated by hBN, we retrieve fundamental material properties that complement and correct the existing data. The method is simple and easy to use for other experiments to retrieve fundamental material properties.
Acknowledgments
This research is funded by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under Grant No. 103.01-2023.138 and carried out by the high-performance cluster at Ho Chi Minh City University of Education, Vietnam.
V.-H.L., D.-N.Ly, and D.-N.Le conceptualized the work and performed analytical formulation; D.-N.Ly, D.-K.D.L., and V.-H.L. carried out numerical calculations; V.-H.L., D.-N.Ly, and D.-N.Le analyzed data, V.-H.L., D.-N.Ly, and D.-N.Le wrote and edited the manuscript.